# Interactions across multiple games: cooperation, corruption, and organizational design

### v2

@@ 5.2 Designer-Optimal Assignments
In this section, we first ask when the designer benefits from the added flexibility of dynamically, reactively assigning teams to tasks. Then, for the interesting case—in which the temptation to deviate on bad games is sufficiently low—we show that optimal reactive team structures take a particular form. We characterize optimal structures both for situations in which the designer can condition task assignments on the occurrence of games as well as when they cannot. Lastly, we consider how the results change as we take the length of a time period to zero.
Under which conditions do dynamically and reactively assigning teams to tasks increase the optimal social value? If full specialization leads to optimal cooperation (Observation 1), then there is no room to improve the social value further. At the other extreme, if there is no share of bad tasks that ensures cooperation, then cooperation on good games is never possible, even with a dynamic assignment. Thus, we focus on situations in which partial specialization leads to optimal outcomes and cooperation on good games can be incentivized, as illustrated in Figure 2.
@@
  • and $\nu^{g}(t_{g})=1$ and $r_{\nu^{g}}$ is sufficiently high to preclude cooperation.
@@
Given the premise of the proposition and $d_{G}\neq d_{B}$, players have a strict incentive to cooperate on one of the games, say game $g$, and are indifferent between cooperating or not on the other, say game $g^{\prime}$. The designer can then make future bad games, those that provide relatively stronger incentives for current cooperation, more likely following game $g^{\prime}$ and less likely following game $g$, keeping the unconditional distribution of games (and tasks) unchanged. As a result, players are now strictly incentivized to cooperate on both games, and the designer can reduce the share of bad tasks in that assignment, as in Example 1, improving the social value. [^13] The converse only holds for the case when the designer does not observe the arrival of games, but not otherwise.
Let us now turn to characterizing designer-optimal reactive team structures. Our first result in this regard establishes that there are essentially two types of teams: those who always cooperate and those who never cooperate.
@@
For intuition, suppose a team were to cooperate only sometimes. The designer could split such a team up into one that only transitions across the states with cooperation in the initial assignment and another that is assigned to the remaining states. Since there would be no periods without cooperation, the first team is now strictly incentivized to cooperate, and so the designer can reduce the share of bad games with cooperation, and Lemma 3 follows. Thus, organizations optimally minimize the share of undesirable cooperation needed to sustain desirable cooperation by keeping the cooperating teams together and eliminating states without cooperation, allowing the incentives coming from future undesirable cooperation to have their maximum impact.
@@ 5.2.1 Good Games Are Observable
Now, we fully characterize the optimal reactive team structures, and we begin with the case in which the designer observes whether a good game occurs.
@@
To understand why observability of the good game’s arrival matters, let us revisit Example 1. In that example, the reactive assignment was a good task followed by a bad task, followed by a bad task with some probability before repeating. The designer can further improve by conditioning on whether a good game actually arrives under the good task. If none arrives, then there is no reason to transition to the bad task, as no incentives are needed. In contrast, if one arrives, then the bad task assignment is needed. Thus, the designer can improve by waiting to transition to the bad task until after the good game actually arrives.
@@
For this analysis, we specialize to the interesting case where the incentive for cooperation in the good game is the one that is binding. In particular, we assume that the temptation to deviate on a current bad game is sufficiently low, specifically that facing a stream of future good games alone would suffice to ensure cooperation on current bad games, i.e., $d_{B}<\frac{\delta}{1-\delta}a_{G}c_{G}$. When this condition holds, a team would cooperate in equilibrium in the bad game, regardless of which games it sees in the future (presuming cooperation). Given this assumption, the binding incentive constraint is on good games, and then the ensuing proposition provides a full characterization of designer-optimal team structures.
@@
By Lemma 3, there are only fully cooperative or fully non-cooperative teams. Proposition 4 below additionally says that if a fully cooperative team exists, it is assigned to the good task until a good game arrives; then, it is assigned to the bad task for a number of periods before returning to the good task. [^14] While it may take an arbitrary number of periods to transition from the good task to the bad task, a reactive assignment conditioning on a finite number of past tasks and games suffices as only whether a good game arrived in the last period matters. The duration of the assignment to the bad tasks is minimized subject to providing enough incentives to cooperate on the good game. If the resulting steady-state distribution of good and bad games with cooperation leads to a net positive social value, the mass of teams in that reactive assignment is maximized, and the teams that never cooperate are assigned to the remaining task to maintain the overall condition on task coverage. For example, when relatively more bad tasks are assigned in the steady-state for cooperating teams than must be assigned to the organization as a whole, the non-cooperative teams would be assigned to the good task. As a result, the non-cooperative teams will be extremal and all assigned to the same task.
@@
Because reactive assignments allow arbitrary (random) conditioning on full histories, the policy class is extremely rich. Our characterization pins down the global optimum within this high-dimensional space and shows that optimality is attained by a simple policy with only two building blocks (fully cooperative vs fully non-cooperative teams) and a minimal bad-block after good arrivals.
@@
Figure 3 illustrates the proposition, showing a reactive assignment in which players always cooperate, and that is part of a designer-optimal reactive team structure. Consider a team currently assigned to the good task. If no game arrives ($g=\emptyset$), it is again assigned to the good task; if a good game arrives ($g=G$), it is instead assigned to the bad task for $N_{B}$ periods. In the next period, with probability $x$, the team is assigned to another bad task before returning to the good task; with complimentary probability, the team is directly assigned to the good task. [^15] For a specific example, let $\delta=\frac{3}{5},c_{G}=c_{B}=1,a_{G}=\frac{1}{2},a_{B}=\frac{3}{4},d_{G}=1$, and $d_{B}=\frac{1}{2}$ (that is, we continue with the parameters in footnote 9 in the context of Example 1). Then, $N_{B}=1$ and $x=\frac{5}{9}$.
[…]
Figure 3: Teams are assigned to task $t_{G}$ with probability 1 until a good game arrives; then, teams are assigned to task $t_{B}$ for $N_{B}$ periods, followed by a random transition either to task $t_{G}$ or to one more task $t_{B}$ before returning to task $t_{G}$. Note that the assignment following a good task conditions on the arrival of a game, whereas the assignment following a bad task does not.
@@ 5.2.2 Good Games Are Unobservable
Again, revisiting Example 1, the improvement we noted at the beginning of Section 5.2.1 in which the designer waited for the arrival of a good game before transitioning is no longer possible. Instead, the designer simply transitions teams directly after the good task, as the incentives are needed if a good game did indeed arrive. As a result, this case does strictly worse than the case in which the arrival of the game is observable (and arrives with probability less than one), but still strictly improves over the case with a static assignment.
@@
Figure 4 illustrates the proposition, showing a reactive assignment with cooperation that is part of a designer-optimal reactive team structure. Note that unlike in the reactive assignment displayed in Figure 3, a team assigned to the good task is assigned to a bad task in the subsequent period (provided $N_{B}>0$) regardless of whether a game arrives or not. [^16] Continuing with the parameter values from footnote 9, we now have $N_{B}=1$ and $x=\frac{5}{27}$. Furthermore, the share of time spent on bad games is $\frac{32}{59}$, whereas it is $\frac{7}{16}$ when the designer observes the arrival of good games.
[…]
Figure 4: Teams are assigned to task $t_{G}$ with probability 1 for one period; then, teams are assigned to task $t_{B}$ for $N_{B}$ periods, followed by a random transition either to task $t_{G}$ or to one more task $t_{B}$ before returning to task $t_{G}$. Note that neither the assignment following a bad task nor a good task conditions on the arrival of a game.
@@
5.2.3 The Continuous Time Limit
We conclude our discussion on designer-optimal reactive team structures by considering how they depend on the length of a time period. In particular, we consider the continuous limit of our model; i.e., when time periods become infinitesimally small and, with them, the probability of a game occurring and the potential frequency of reassignments. We will see that whether the designer can improve upon static assignments as the period length shrinks depends on whether the arrival of good games is observable. This suggests that dynamic assignments in the unobservable case are only useful in situations where at most one good game arrives over a nontrivial span of time.
So far, a period has been of length $1$. Suppose now that a period is of length $T<1$, which implies two changes. First, players in period $\tau$ discount payoffs received in period $\tau^{\prime}>\tau$ by $\delta^{(\tau^{\prime}-\tau)T}$. Second, the probability of arrivals of good and bad games on the two tasks are now $a_{G}T$ and $a_{B}T$, respectively; and, as $T\rightarrow 0$, games occur according to a Poisson process. As before, the designer can change assignments dynamically and reactively in each period. [^17] There might be institutional constraints that prevent the designer from adjusting assignments every period. We do not study this variant but do not believe it will change the basic intuition of the results as it just adds constraints to the designer.
We first note that as $T$ decreases, while the designer’s payoff from any fixed static team structure stays the same, the designer’s payoff from an optimal static team structure increases. The expected arrival rate of games is unchanged, but the timing of arrivals is more continuous rather than concentrated at the end of a full period. Under exponential discounting, a continuous stream of possible arrivals within, say, the next two periods is valued more highly than a single arrival exactly one period away. Thus, the same expected rate of future cooperation is effectively worth more. As a result, the designer can sustain cooperation with a smaller share of bad tasks, and the social value under the optimal static assignment increases as $T\to 0$.
Now consider reactive team structures and the case where the designer observes good games occurring (Section 5.2.1). The optimal reactive team structure does not change substantively: Consider a reactive assignment in which players always cooperate, which is part of an optimal reactive team structure. In the limit as $T\rightarrow 0$, following a good game, teams spend some fixed amount of time on the bad task before returning to the good task, where they wait until a good game arrives, as before. Furthermore, as was true in the $T=1$ case, the social value is higher when reactive assignments are possible compared to static assignments, as the designer can wait for the arrival of a good game. [^18] Thus, combining this conclusion with the observation in the previous paragraph, both the reactive and static assignments are improving, but the reactive is strictly better.
Consider next an optimal reactive team structure when the designer does not observe good games occurring (Section 5.2.2) and a reactive assignment in which players always cooperate. As $T\rightarrow 0$, players are immediately transitioned out of the good task, and the time spent on ensuing bad tasks also becomes arbitrarily short. In other words, players are randomly assigned to the good or bad tasks with time-independent probabilities, and the designer cannot benefit from a reactive assignment. Intuitively, if time periods are long, then there is a high chance that a team on the good task faced a good game. Thus, it is as if the designer knows a good game took place. When the length of a period shrinks, the designer loses this knowledge. As a result, making bad games relatively more likely following a good game becomes infeasible. This implies that the gains the designer has from reactive assignments compared to static assignments when good games are not observed vanishes as the length of the period goes to zero and arbitrarily frequent adjustments are possible. [^19] One can show that the designer’s payoff improves as $T\rightarrow 0$. The designer’s payoff from a static assignment is also improving as $T\rightarrow 0$, and it is the relative improvement that vanishes. Thus, whether there is a non-vanishing improvement over the static assignment as the time period shrinks depends on the observability of the arrival of good games.
@@ 6 Concluding Remarks
A systemic approach is important in understanding and designing organizations. Nominally independent tasks and situations become dependent if team members interact repeatedly. Our analysis suggests that organizations can overcome some of the undesirable spillover effects through intentional design. For example, the Mollen Commission (1994), discussing corruption and policing, describes an empirical regularity showing the importance of a systemic view: “[the code of silence] is strongest where corruption is most frequent. This is because the loyalty ethic is particularly powerful in crime-ridden precincts where officers most depend upon each other for their safety each day.” The “loyalty ethic” corresponds to cooperation in bad situations and the “depend[ence] upon each other for their safety” describes the benefits from cooperation in good situations. The report then suggests that “crime-ridden precincts” are those with a high frequency of the good game (e.g., more dangerous situations), which facilitates cooperation in bad games—an important instantiation of spillover effects. Our results on reshuffling, specialization and dynamic assignments suggest organizational design can be used to address these concerns. The police force can specialize teams that encounter dangerous situations on the job (who then also accept bribes) and thereby reduce the number of officers in such situations to minimize corruption. The police force can optimize further by rotating such teams to tasks that foster teamwork, building up a norm of cooperation before heading back “into the field.”
Our analysis also relates to the cohesion of communities and societies. Banfield (1958); Putnam (2000), in their studies of an Italian city and the US over time, respectively, document these spillovers. Banfield (1958) describes communities with nuclear families as uncooperative actors. As Lemma 1 shows, a lack of interaction in some dimensions may lead to a breakdown of all cooperation (public goods, commons, etc.). Similarly, the resulting need for cooperation within the family can create powerful incentives to act as a single unit, which could enhance the ability to coordinate on being corrupt. Relatedly, Putnam (2000) warns about the consequences of an erosion of civic interactions (“good cooperation”) in the US for other kinds of institutions that require cooperation, e.g., political discourse. Lemma 1 supports his warning of spillovers from cooperation on leisurely activities, such as bowling, to more general foundations of civic behavior. It further provides some insight into what may have caused such a decline in civic norms: formal institutions, such as insurance markets and markets, replacing informal interactions—such as reciprocal insurance and favor exchange—thereby reducing necessary cooperation.
@@ 7 Bells and Whistles
In closing, we discuss a few extensions.
@@
When assigning teams to tasks, we focused on tasks that either had only good game or only bad games. Hybrid tasks that involve some random game arrivals could be accommodated, albeit with complications in notation and cases, but would yield similar results throughout. Also, in our model, the set of tasks is fixed. However, when supporting optimal cooperation is challenging, it can be advantageous to introduce new tasks that provide additional benefits from cooperation. These new tasks might not benefit society or directly increase organizational productivity, but they might help sustain cooperation in good games. For instance, scheduling “team-building” exercises, social events, and non-production related contests in which team members must cooperate with each other can help induce cooperation in other games as well. Many organizations use such team-building to enhance the loyalty that people feel towards the units in which they are involved, fostering cooperation. Although such events may be time-consuming and thus reduce the time that teams spend directly on productive activities, they may end up enhancing productivity and substitute for bad games by enabling cooperation in good games when that might have otherwise been infeasible.
@@
Appendix A Comparative Statics in Payoffs and Design Implications
[…]
(a) Initially: $d_{B}>d_{G}$ but $c_{G}$ is too low for optimal cooperation.
(b) Increase: $c_{G}$ to $c_{G}^{\prime}$
[…]
(c) Initially: $d_{G}>d_{B}$.
(d) Decrease: $d_{G}$ to $d_{G}^{\prime}$.
Appendix B Proofs

### v3

@@ 5.2 Designer-Optimal Assignments
Throughout what follows, we focus on the most relevant case in which the incentive to deviate in bad games, $d_{B}$, is sufficiently low so that the incentives provided to cooperate in good tasks are the binding constraint. [^13] When, instead, $d_{G}$ is sufficiently low, full specialization ensures optimal cooperation (Observation 1).
In this section, we first ask when the designer benefits from the added flexibility of dynamically and reactively assigning teams to tasks. We then show that optimal reactive team structures take a particular form. We characterize these structures both when the designer can condition task assignments on the occurrence of games and when they cannot. In an optimal team structure, teams that cooperate are assigned to the good task until a good game arrives, at which point they are moved to the bad task for some amount of time before returning to the good task. Teams that do not cooperate pick up the slack to satisfy the exogenous task constraint. Finally, we examine how our results change as the length of a time period shrinks to zero.
When does dynamically and reactively assigning teams to tasks increase the optimal social value? If full specialization leads to optimal cooperation (Observation 1), then there is no room to improve the social value further. At the other extreme, if there is no share of bad tasks that ensures cooperation, then cooperation on good games is never possible, even with a dynamic assignment. Thus, we focus on situations in which partial specialization leads to optimal outcomes and cooperation on good games can be incentivized, as was illustrated in Figure 2. In our running example, a further improvement to the static assignment $\nu^{coop}$ was possible as $d_{G}<d_{B}$. The next proposition shows that the general condition is unequal deviation payoffs, $d_{G}\neq d_{B}$.
@@
  • and $\nu^{g}(t_{g})=1$ and $r_{g}$ is sufficiently high to preclude cooperation if $g=B$.
@@
Given the premise of the proposition and $d_{G}\neq d_{B}$, players have a strict incentive to cooperate on one of the games, say game $g$, and are indifferent between cooperating or not on the other, say game $g^{\prime}$. The designer can then make future bad games, those that provide relatively stronger incentives for current cooperation, more likely following game $g^{\prime}$ and less likely following game $g$, keeping the unconditional distribution of games (and tasks) unchanged. As a result, players are now strictly incentivized to cooperate on both games, allowing the designer to reduce the share of bad tasks in that assignment, as in the running example, and thereby improve the social value. [^14] The converse of the proposition statement also holds, but only for the case when the designer does not observe the arrival of games.
We now turn to characterizing designer-optimal reactive team structures.
Our first result makes it easier to describe the optimal structure as it establishes that there are essentially two types of teams: those that always cooperate and those that never cooperate.
@@
For intuition, suppose a type of assignment in the support of $\lambda$ induces cooperation only sometimes. The designer could split such an assignment into two assignments: one that transitions only across the states with cooperation in the initial assignment and another that is assigned to the remaining states. Since there would be no periods without cooperation, teams with the first assignment are now strictly incentivized to cooperate, allowing the designer to reduce the share of bad games with cooperation, and Lemma 3 follows. Thus, organizations optimally minimize the share of undesirable cooperation needed to sustain desirable cooperation by keeping cooperating teams together and eliminating states without cooperation, thereby allowing the incentives arising from future undesirable cooperation to have their maximum impact.
@@ 5.2.1 Good Games Are Observable
Now, we fully characterize the optimal reactive team structures, beginning with the case in which the designer observes whether a good game occurs.
@@
To understand why observability of the good game’s arrival matters, let us revisit the dynamic assignment in which a good task is followed by a bad task, followed by a bad task with some probability before repeating. The designer can further improve by conditioning on whether a good game actually arrives under the good task. If none arrive, then there is no reason to transition to the bad task, as no incentives are needed. In contrast, if one arrives, then the bad task assignment is needed. Thus, the designer can improve by waiting to transition to the bad task until after a good game actually arrives.
@@
For this analysis, we specialize to the interesting case in which the incentive for cooperation in the good game is the binding constraint. In particular, we assume that the temptation to deviate on a current bad game is sufficiently low, specifically that facing a stream of future good games alone would suffice to ensure cooperation on current bad games, i.e., $d_{B}<\frac{\delta}{1-\delta}a_{G}c_{G}$. When this condition holds, a team would cooperate in equilibrium in the bad game, regardless of which games it sees in the future (presuming cooperation). Given this assumption, the binding incentive constraint is on good games, and the ensuing proposition provides a full characterization of designer-optimal team structures.
@@
By Lemma 3, there are only fully cooperative or fully non-cooperative teams. Proposition 4 below additionally says that if a fully cooperative team exists, it is assigned to the good task until a good game arrives; then, it is assigned to the bad task for a number of periods before returning to the good task. [^15] While it may take an arbitrary number of periods to transition from the good task to the bad task, a reactive assignment conditioning on a finite number of past tasks and games suffices, as only whether a good game arrived in the last period matters. The duration of the assignment to the bad tasks is minimized, subject to providing enough incentives to cooperate on the good game. If the resulting steady-state distribution of good and bad games with cooperation leads to a net positive social value, the mass of teams in that reactive assignment is maximized, and the teams that never cooperate are assigned to the remaining task to maintain the overall condition on task coverage, as before.
@@
Because reactive assignments allow arbitrary (random) conditioning on whatever the designer observes, the policy class can be extremely rich. Our characterization pins down the global optimum in this high-dimensional space. It shows that optimality is attained by a simple policy with only two building blocks (fully cooperative vs. fully non-cooperative teams) and a minimal sequence of bad tasks after good arrivals.
@@
To illustrate the proposition, consider again Example 1. The proposition implies that the reactive assignment in which players always cooperate, and that forms part of a designer-optimal reactive team structure, can be described by the following Markov chain
$\bordermatrix{&(t_{G},G)&(t_{G},\emptyset)&(t_{B},B)_{1}&(t_{B},\emptyset)_{1}&(t_{B},B)_{2}&(t_{B},\emptyset)_{2}\cr(t_{G},G)&0&0&a_{B}&1-a_{B}&0&0\cr(t_{G},\emptyset)&a_{G}&1-a_{G}&0&0&0&0\cr(t_{B},B)_{1}&(1-x)a_{G}&(1-x)(1-a_{G})&0&0&xa_{B}&x(1-a_{B})\cr(t_{B},\emptyset)_{1}&(1-x)a_{G}&(1-x)(1-a_{G})&0&0&xa_{B}&x(1-a_{B})\cr(t_{B},B)_{2}&a_{G}&1-a_{G}&0&0&0&0\cr(t_{B},\emptyset)_{2}&a_{G}&1-a_{G}&0&0&0&0\cr}.$
Specifically, consider a team currently assigned to the good task. If no game arrives, the state is $(t_{G},\emptyset)$, and it is assigned again to the good task. Conditional on this transition, a good game arrives with probability $a_{G}$, in which case the state becomes $(t_{G},G)$; otherwise, the state remains $(t_{G},\emptyset)$. If the state is $(t_{G},G)$, the team is then assigned to the bad task for at least one period. In particular, if a bad game arrives, the next state is $(t_{B},B)_{1}$, and if no bad game arrives, the next state is $(t_{B},\emptyset)_{1}$. As in the transition matrix (6), the team may then be assigned to another bad task with probability $x$, or transition back to the good task. For the parameter values in Example 1, one can show that, indeed, for $N_{B}=1$ and $x=5/9$, players are indifferent between cooperating and deviating when facing a good game, as optimality requires.
@@ 5.2.2 Good Games Are Unobservable
Again, revisiting Example 1, the improvement we noted at the beginning of Section 5.2.1 in which the designer waited for the arrival of a good game before transitioning is no longer possible. Instead, the designer simply transitions teams directly after the good task, as the incentives are needed if a good game did indeed arrive; that is, the teams transition according to the matrix (6). As a result, this case does strictly worse than the case in which the arrival of the game is observable (and arrives with probability less than one), but still strictly improves over the case with a static assignment.
@@
The proposition confirms that the Markov Chain transition matrix (6) characterizes the structure of an optimal reactive assignment when cooperation is optimal and good games are unobservable. Given the parameters in Example 1, we obtain $N_{B}=1$ and $x=5/27$. In the corresponding steady state, fully cooperative teams spend $32/59$ of their time on bad tasks, compared to $7/16$ when the designer observes the arrival of good games. Thus, social value is strictly lower in the unobservable case (which still is an improvement relative to the static case in which the share spent on bad tasks is $\nu^{coop}(t_{B})=2/3$).
@@
In both the observable and unobservable cases, optimal reactive team structures have a simple form: they rely on only two building blocks—fully cooperative and fully non-cooperative teams—and feature a finite sequence of bad-task assignments following good-task assignments. The key difference is that when good games are observable, the transition to the bad task can be conditioned on a good game’s actual arrival, so it is not forced when it is not needed, whereas when the arrival is unobservable, the designer must transition regardless. Consequently, the observable case achieves the same incentive provision with fewer bad tasks in steady state and strictly higher social value.
@@ 6 Concluding Remarks
A systemic approach is important in understanding and designing organizations. Nominally independent tasks and situations become dependent if team members interact repeatedly. Our analysis suggests that organizations can overcome some of the undesirable spillover effects through intentional design. For example, the Mollen Commission (1994), discussing corruption and policing, describes an empirical regularity showing the importance of a systemic view: “[the code of silence] is strongest where corruption is most frequent. This is because the loyalty ethic is particularly powerful in crime-ridden precincts where officers most depend upon each other for their safety each day.” The “loyalty ethic” corresponds to cooperation in bad situations, and the “depend[ence] upon each other for their safety” describes the benefits from cooperation in good situations. The report then suggests that “crime-ridden precincts” are those with a high frequency of the good game (e.g., more dangerous situations), which facilitates cooperation in bad games—an important instantiation of spillover effects. Our results on reshuffling, specialization, and dynamic assignments suggest organizational design can be used to address these concerns. The police force can specialize teams that encounter dangerous situations on the job (who then also accept bribes) and thereby reduce the number of officers in such situations to minimize corruption. The police force can optimize further by rotating such teams to tasks that foster teamwork, building up a norm of cooperation before heading back “into the field.”
Our analysis also relates to the cohesion of communities and societies. Banfield (1958); Putnam (2000), in their studies of an Italian city and the US over time, respectively, document these spillovers. Banfield (1958) describes communities with nuclear families as uncooperative actors. As Lemma 1 shows, a lack of interaction in some dimensions may lead to a breakdown of all cooperation (public goods, commons, etc.). Similarly, the resulting need for cooperation within the family can create powerful incentives to act as a single unit, which could enhance the ability to coordinate on being corrupt. Relatedly, Putnam (2000) warns about the consequences of the erosion of civic interactions (“good cooperation”) in the US for other kinds of institutions that require cooperation, e.g., political discourse. Lemma 1 supports his warning of spillovers from cooperation on leisurely activities, such as bowling, to more general foundations of civic behavior. It further provides some insight into what may have caused such a decline in civic norms: formal institutions, such as insurance markets and markets, replacing informal interactions—such as reciprocal insurance and favor exchange—thereby reducing necessary cooperation.
@@ 7 Bells and Whistles
In closing, we discuss a few extensions and other variants of the model.
@@
When assigning teams to tasks, we focused on tasks that either had only good games or only bad games. Hybrid tasks that involve some random game arrivals could be accommodated, albeit with complications in notation and cases, but would yield similar results throughout. Also, in our model, the set of tasks is fixed. However, when supporting optimal cooperation is challenging, it can be advantageous to introduce new tasks that provide additional benefits from cooperation. These new tasks might not benefit society or directly increase organizational productivity, but they might help sustain cooperation in good games. For instance, scheduling “team-building” exercises, social events, and non-production related contests in which team members must cooperate with each other can help induce cooperation in other games as well. Many organizations use such team-building to enhance the loyalty that people feel towards the units in which they are involved, fostering cooperation. Although such events may be time-consuming and thus reduce the time that teams spend directly on productive activities, they may end up enhancing productivity and substitute for bad games by enabling cooperation in good games when that might have otherwise been infeasible.
@@
The continuous time limit.
Time periods in our model are discrete, and a natural question to ask is how the previous results depend on the length of a time period. To this end, we consider the continuous limit of our model; i.e., when time periods become infinitesimally small and, with them, the probability of a game occurring and the potential frequency of reassignments. We will see that whether the designer can improve upon static assignments as the period length shrinks with a dynamic assignment depends on whether the arrival of good games is observable. This suggests that dynamic assignments in the unobservable case are only useful in situations where at most one good game arrives over a nontrivial span of time.
So far, a period has been of length $1$. Suppose now that a period is of length $T<1$, which implies two changes. First, players in period $\tau$ discount payoffs received in period $\tau^{\prime}>\tau$ by less, specifically, by $\delta^{(\tau^{\prime}-\tau)T}$. Second, the probabilities of arrival of good and bad games on the two tasks are also reduced, specifically to $a_{G}T$ and $a_{B}T$, respectively; and, as $T\rightarrow 0$, games occur according to a Poisson process. As before, the designer can change assignments dynamically and reactively in each period. [^16] There might be institutional constraints that prevent the designer from adjusting assignments every period. We do not study this variant, but we do not believe it will change the basic intuition of the results, as it just adds constraints to the designer.
We first note that as $T$ decreases, while the designer’s payoff from any fixed static team structure stays the same, the designer’s payoff from an optimal static team structure increases. The expected arrival rate of games remains unchanged, but the timing of arrivals is more continuous rather than concentrated at the end of a full period. Under exponential discounting, a continuous stream of possible arrivals within, say, the next two periods is valued more highly than a single arrival exactly one period away. Thus, the same expected rate of future cooperation is effectively worth more. As a result, the designer can sustain cooperation with a smaller share of bad tasks, and the social value under the optimal static assignment increases as $T\to 0$.
Now consider reactive team structures and the case where the designer observes good games occurring (Section 5.2.1). The optimal reactive team structure does not change substantively: consider a reactive assignment in which players always cooperate, which is part of an optimal reactive team structure. In the limit as $T\rightarrow 0$, following a good game, teams spend some fixed amount of time on the bad task before returning to the good task, where they wait until a good game arrives, as before. Furthermore, as was true in the $T=1$ case, the social value is higher when reactive assignments are possible compared to static assignments, as the designer can wait for the arrival of a good game. [^17] Thus, combining this conclusion with the observation in the previous paragraph, both the reactive and static assignments are improving, but the reactive assignment remains a strict improvement.
Consider next an optimal reactive team structure when the designer does not observe good games occurring (Section 5.2.2) and a reactive assignment in which players always cooperate. As $T\rightarrow 0$, players are immediately transitioned out of the good task, and the time spent on ensuing bad tasks also becomes arbitrarily short. In other words, players are randomly assigned to the good or bad tasks with time-independent probabilities, and the designer cannot benefit from a reactive assignment. Intuitively, if time periods are long, then there is a high chance that a team on the good task faced a good game. Thus, it is as if the designer knows a good game took place. When the length of a period shrinks, the designer loses this knowledge. As a result, making bad games relatively more likely following a good game becomes infeasible. This implies that the gains the designer has from reactive assignments compared to static assignments when good games are not observed vanish as the length of the period goes to zero and arbitrarily frequent adjustments are possible. [^18] One can show that the designer’s payoff improves as $T\rightarrow 0$. The designer’s payoff from a static assignment also improves as $T\rightarrow 0$, and it is the relative improvement that vanishes. Thus, whether there is a non-vanishing improvement over the static assignment as the time period shrinks depends on the observability of the arrival of good games.
Appendix A Proofs